OpenStudy (anonymous):

Will give Medal to #1 answer! Will give fan for everyone else!

OpenStudy (anonymous):

How do you find the square root of a decimal?

OpenStudy (anonymous):

@razor99 @Lady.Liv1776 @Loser66 @leonardo0430 @Legends @ParthKohli @Omgthatgirl @austinL @sammixboo

OpenStudy (anonymous):

you want to convert it to a rational number. Example: √1.44 = √144/√100 = 12/10 = 1.2

OpenStudy (anonymous):

Sorry, here's the link: http://www.ask-math.com/square-root-of-decimals.html

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

Anytime :)

OpenStudy (anonymous):

Does anyone know how to find: |x - 2| - 4|-6| where x<2